# RoomPainter: View-Integrated Diffusion for Consistent Indoor Scene Texturing

### v1

@@ 3.4 Related View-based Attention
where $\tilde{K}_{n}=\begin{bmatrix}K_{1}\\ K_{2}\\ \vdots\\ K_{R}\end{bmatrix}$ and $\tilde{V}_{n}=\begin{bmatrix}V_{1}\\ V_{2}\\ \vdots\\ V_{R}\end{bmatrix}$. In the first room-scale texturing stage, we consider the related viewpoints $R$ for view $n$ as its adjacent views, i.e., its left and right view. When sampling at instance scale, we define the related views as all the $N$ predefined camera view.
@@ 4 Experiments
Figure 4: Qualitative comparisons. Text2Tex-H [5] suffers from occlusion and visible seams. Text2Tex-C [5] struggles to maintain style consistency across all instances. SceneTex [6] produces unrealistic textures and results in blurry regions. In contrast, our method generates high-quality textures while preserving overall style consistency across instances in the scene. Ceilings and back-facing walls are excluded for improved visualizations.
@@ 4.1 Baseline Implementation Details
  • Text2Tex-C [5]: For comparative purposes, we employ the original configuration of Text2Tex, wherein each room instance, including the interior walls and furniture, is textured using an instance-specific text prompt. And for interior walls texturing, we utilize the text prompt such as: ”A Chinese style living room, no furniture” and sampled 12 perspectives same as Text2Tex-H. All the textured instances are subsequently combined, forming the output textured room.
@@ 4.2 Implementation Details
In this experiment, we used the SDXL model [28] as the text-to-image (T2I) generator and incorporated depth information through ControlNet [51, 1]. For the text prompt, we implemented view-specific prompt combining the holistic prompt with the names of the types of objects present within the field of view. For sampling, we applied the DDPM [16] as the sampler, setting the sampling steps to 50. We used PyTorch3D [31] for rendering and texture projection. Additionally, UV atlas generation for the mesh was handled with Xatlas [46]. All experiments were performed on an NVIDIA RTX A6000 GPU. Further implementation details are available in the supplementary materials.
@@ 4.3 Quantitative Analysis
Our quantitative experiment were conducted across 10 scenes including 5 bedroom and 5 living room from the 3D-FRONT dataset [12], using five different text styles totally. To evaluate the generated texture, we calculated both the CLIP Score(CLIP) [15], which assesses the alignment between generated texture and the textual description, and the CLIP Aesthetic Score(AS) [35], which measures texture quality. We also compared the average processing time of each texture generation method in minutes. The results are summarized in Table.1. We render 10 images from novel views for each room to calculate the average of CLIP and AS. Due to the presence of numerous distinct objects within the room, the methods in the baselines do not adjust the text prompts for texture generation based on the variation of objects from different viewpoints. Therefore, we employ a holistic prompt to compute the CLIP score. Since our evaluation metric involves image sampling from new viewpoints, Text2Tex-H, which fails to address the occlusion problem, generates textures containing large occluded regions, resulting in the lowest performance in terms of the metric. However, due to the limited number of sampled viewpoints, it exhibits the fastest generation speed. We configure SceneTex according to its original setup, sampling from a sphere centered at the room’s center. As a result, regions not observed by SceneTex often exhibit blurriness. Meanwhile, in addition to optimizing the texture field using VSD, SceneTex also requires training LoRA, which results in the longest processing time. Text2Tex-C employs an instance-specific approach for texture generation, utilizing different prompts for each object, which results in higher performance metrics. Additionally, it incorporates viewpoint selection and texture optimization processes, leading to a longer computation time.
@@ 4.4 Qualitative Analysis
We provide a visual comparison in fig.4. Specifically, Text2Tex-H [5] shows that simply applying object texture generation method to indoor scene texturing task cannot resolve the occlusion problem (see the large black areas in the first column of fig.4) and contain obvious seams. Text2Tex-C [5] texturing the room completely but suffering from obvious seams and global style consistency across objects. Although SceneTex introduces the use of cross-attention to generate textures for occluded regions, noticeable blurriness still persists in these areas. Furthermore, due to its optimization-based approach, SceneTex often produces unrealistic textures. In contrast, our method generates globally consistent textures with a unified style in the first stage. In the second stage, through per-instance inpainting and refinement, we successfully maintain global style consistency while addressing issues of occlusion and unrealistic textures.
@@ 4.5 Ablation Studies
Effectiveness of the MVIS. We employ a two-stage approach to generate textures for the room. In the first stage, we introduce MVIS to enhance the global consistency of the textures. Qualitative results from this stage are shown in second column of the Fig. 5. The figures demonstrate that without MVIS, texture consistency across multiple viewpoints can be significantly impact. We further conduct quantitative ablation on dataset adopted in previous experiments. Results are shown in the seventh row of Tab. 1. The quantitative experimental results clearly demonstrate that MVIS improves the quality of the generated textures. Effectiveness of the Related View-based Attention. We incorporate the Related View-based Attention module in both stages of texture generation to enhance consistency. Both qualitative(the first column of Fig. 5) and quantitative experiments(the fifth row of the Tab. 1) demonstrate that using only MVIS or MVRS is insufficient to generate consistent and high-quality textures. The Related View-based Attention module effectively enhances consistency while ensuring the quality of the generated textures. Effectiveness of the MVRS. In the second stage of texture generation, we use MVRS to generate textures for the occluded regions of each instance, while also refining the textures generated in the first stage. Quantitative experimental results (seventh row in Tab. 1) show that without MVRS, the quality of the textures significantly deteriorates. This is because, without MVRS, the textures generated in the first stage suffer from low-quality results due to occlusion. We will provide additional qualitative materials demonstrating the effectiveness of MVRS in the supplementary materials.
@@ 5 Conclusion
We propose an indoor scene texturing framework named RoomPainter that generates high-quality textures with multi-view consistency and time efficiency. By leveraging the carefully designed modules including view-weighted texture rectification and related view-based attention, proposed view-integrated diffusion preserves the cross-view consistency during the diffusion process. Qualitative and quantitative results demonstrate the effectiveness of RoomPainter in indoor scene texturing, especially in global consistency and generation efficiency.

### v2

@@ 3.4 Related View-based Attention
where $\tilde{K}_{n}=\begin{bmatrix}K_{1}\ K_{2}\ \ldots\ K_{R}\end{bmatrix}^{T}$ and $\tilde{V}_{n}=\begin{bmatrix}V_{1}\ V_{2}\ \ldots\ V_{R}\end{bmatrix}^{T}$. In the first room-scale texturing stage, we consider the $R$ related viewpoints for view $n$ as its adjacent views, i.e., its left and right view. When sampling at instance scale, we define the related views as all the $N$ predefined camera view.
@@ 4 Experiments
Figure 4: Qualitative comparisons. Text2Tex-H [6] suffers from occlusion and visible seams. Text2Tex-C [6] struggles to maintain style consistency across all instances. SceneTex [7] produces unrealistic texture and results in blurry regions. In contrast, our method generates high-quality texture while preserving overall style consistency across instances in the scene. Ceilings and back-facing walls are excluded for improved visualizations.
@@ 4.1 Baseline Implementation Details
  • Text2Tex-C [6]: For comparative purposes, we employ the original configuration of Text2Tex, wherein each room instance, including the interior walls and furniture, is textured using an instance-specific text prompt. And for interior walls texturing, we utilize the text prompt such as: "A Chinese style living room, no furniture" and sampled 12 perspectives same as Text2Tex-H. All the textured instances are subsequently combined, forming the output textured room.
@@ 4.2 Implementation Details
In this experiment, we used the SDXL model [28] as the text-to-image generator and incorporated depth information through ControlNet [56, 1]. For the text prompt, we implemented view-specific prompt combining the holistic prompt with the names of the types of objects present within the field of view. For sampling, we applied the DDPM [17] as the sampler, setting the sampling steps to 50. We used PyTorch3D [31] for rendering and texture projection. Additionally, UV atlas generation for the mesh was handled with Xatlas [48]. All experiments were performed on an NVIDIA RTX A6000 GPU. Further implementation details are available in the supplementary materials.
@@ 4.3 Quantitative Analysis
Our quantitative experiment were conducted across 10 scenes including 5 bedroom and 5 living room from the 3D-FRONT dataset [13], using five different text styles totally. To evaluate the generated texture, we calculated both the CLIP Score(CLIP) [16], which assesses the alignment between generated texture and the textual description, and the CLIP Aesthetic Score(AS) [35], which measures texture quality. We also compared the average processing time of each texture generation method in minutes. The results are summarized in Table.1. We render 10 images from novel views for each room to calculate the average of CLIP and AS. Due to the presence of numerous distinct objects within the room, the methods in baselines do not adjust the text prompts for texture generation based on the variation of objects from different viewpoints. Therefore, we employ a holistic prompt to compute the CLIP score. Since our evaluation metric involves image sampling from new viewpoints, Text2Tex-H, which fails to address the occlusion problem, generates texture containing large occluded regions, resulting in the lowest performance in terms of the metric. However, due to the limited number of sampled viewpoints, it exhibits the fastest generation speed. We configure SceneTex according to its original setup, sampling from a sphere centered at the room’s center. As a result, regions not observed by SceneTex often exhibit blurriness. Meanwhile, in addition to optimizing the texture field using VSD, SceneTex also requires training LoRA, which results in the longest processing time. Text2Tex-C employs an instance-specific approach for texture generation, utilizing different prompts for each object, which results in higher performance metrics. Additionally, it incorporates viewpoint selection and texture optimization processes, leading to a longer computation time.
@@ 4.4 Qualitative Analysis
We provide a visual comparison in fig.4. Specifically, Text2Tex-H [6] shows that simply applying object texture generation method to indoor scene texturing task cannot resolve the occlusion problem (see the large black areas in the first column of fig.4) and contain obvious seams. Text2Tex-C [6] texturing the room completely but suffering from obvious seams and global style consistency across objects. Although SceneTex introduces the use of cross-attention to generate texture for occluded regions, noticeable blurriness still persists in these areas. Furthermore, due to its optimization-based approach, SceneTex often produces unrealistic texture. In contrast, our method generates globally consistent texture with a unified style in the first stage. In the second stage, through per-instance inpainting and refinement, we successfully maintain global style consistency while addressing issues of occlusion and unrealistic texture.
@@ 4.5 Ablation Studies
As shown in Fig.5, all images are raw samples from the final step, thus revealing inherent multi-view inconsistencies.
Effectiveness of the MVIS. In the first stage of texturing, we introduce MVIS to enhance the global consistency of the texture. Qualitative results from this stage are shown in the second column of Fig.5. The figure demonstrates that without MVIS, texture consistency across multiple viewpoints can be significantly impact. We further conduct quantitative ablation on dataset adopted in previous experiments. Results are shown in the seventh row of Tab.1. The quantitative experimental results clearly demonstrate that MVIS improves the quality of the generated texture.
Effectiveness of the Related View-based Attention. We incorporate the Related View-based Attention module in both stages of texture generation to enhance consistency. Both qualitative(the first column of Fig.5) and quantitative experiments(the fifth row of the Tab.1) demonstrate that using only MVIS or MVRS is insufficient to generate consistent and high-quality texture. The Related View-based Attention module effectively enhances consistency while ensuring the quality of the generated texture.
Effectiveness of the MVRS. In the second stage of texturing, we use MVRS to generate texture for the occluded regions of each instance and refine the texture generated in the first stage. Quantitative experimental results (seventh row in Tab.1) show that without MVRS, the quality of the texture significantly deteriorates. This is because, without MVRS, the texture generated in the first stage suffer from low-quality results due to occlusion. We will provide additional qualitative materials demonstrating the effectiveness of MVRS in the supplementary materials.
@@ 5 Conclusion
We propose an indoor scene texturing framework named RoomPainter that generates high-quality texture with multi-view consistency and time efficiency. By leveraging the carefully designed modules including view-weighted texture rectification and related view-based attention, proposed view-integrated diffusion preserves the cross-view consistency during the diffusion process. Qualitative and quantitative results demonstrate the effectiveness of RoomPainter in indoor scene texturing, especially in global consistency and generation efficiency.
Supplementary Material
Appendix A Implementation Details
A.1 Text Prompt of Stage1
In the first stage of texture generation, we use different text prompts when sampling the image corresponding to each view. To determine the text prompt for a specific view, we calculate the proportion of pixels rendered for each instance in the view relative to the total number of pixels. If this proportion exceeds a set threshold $\sigma$, we include the text corresponding to that instance in the global prompt. For example, taking “A Chinese style bedroom” as the holistic prompt, if the proportions of instances: “single bed”, “wardrobe” and “chair” exceed $\sigma$, the text prompt used for sampling this view becomes “A Chinese style bedroom with single bed, wardrobe and chair.” In our implementation, the threshold $\sigma$ is set to $0.01$.
A.2 Text Prompt of Stage2
During the texture generation phase for each instance, we derive the style prompt from the holistic prompt. For example, in “A Chinese style bedroom”, the style prompt is “Chinese”. When applying MVRS to a room without furniture, we use the same prompt for all viewpoints, such as “A Chinese style bedroom, without furniture.” For individual furniture instances, we use view-specific prompts. For example, for the instance “single bed”, the prompt would be “A Chinese-style single bed, [DIR] view” where [DIR] represents the relative position of the viewpoint. The [DIR] value changes based on the azimuth angle of the viewpoint and may include directions such as ‘front’, ‘front side’, ‘rear’, ‘side’, or ‘top-down’.
A.3 Time steps for MVIS
During the MVIS sampling process, we adopt different strategies based on the time step $t$. When $t\in\{T\dots 0.9T\}$, we use the standard diffusion sampling method. For $t\in\{0.9T\dots 0.5T\}$, the MVIS is applied. During $t\in\{0.5T\dots 0.3T\}$, we alternate between diffusion sampling and the MVIS. Finally, for $t\in\{0.3T\dots 0\}$, we revert to the standard diffusion sampling method. This multi-stage sampling strategy accelerates the overall sampling process while maintaining high texture quality.
Method PF $\uparrow$ VQ $\uparrow$ TG $\uparrow$
Text2Tex-H [6] 3.06 2.72 2.43
Text2Tex-C [6] 3.22 3.10 3.15
SceneTex [7] 3.47 3.28 3.42
Ours 4.00 4.13 4.31
Table 2: User study result. We report the prompt fidelity (PF), visual quality (VQ) and texture-geometry alignment (TG) results for user study. We show that our method produces high quality textures that coherent with input text prompt.
A.4 Time steps for MVRS
Similar to the multi-stage sampling strategy in MVIS, we also adopt a multi-stage approach in MVRS to accelerate the process and improve texture quality. When $t\in\{T\dots 0.8T\}$, we use the standard RePaint sampling method without projecting images to texture space. For $t\in\{0.8T\dots 0.5T\}$, we apply the standard MVRS method. During $t\in\{0.5T\dots 0.3T\}$, we alternate between diffusion sampling and the MVIS. Finally, for $t\in\{0.3T\dots 0\}$, we switch back to the standard diffusion sampling method.
A.5 Camera selection of both stage
Our two-phase camera configuration:
Global Phase: N cameras (N=6) are positioned to look at the center of the room, with a radius equal to half the shortest axis of the room.
Instance Phase: N cameras (N=12 for room-frame repainting, N=9 for furniture repainting) are positioned around each instance, looking at its center. The camera distance is set to 0.95 times of the diagonal length of the furniture’s bounding box when repainting furniture. For room-frame repainting, the same camera parameters as global phase are used, but with a field of view (FOV) of 80 degrees.
A.6 Parameters
Throughout all sampling processes, we set the inference steps to $50$. The classifier-free guidance scale is linearly reduced from $10.0$ to $7.0$ over the sampling time steps. The ControlNet conditioning scale is fixed at 1. Additionally, the exp parameter in $dynamic\_merge$ is linearly increased from $1.0$ to $6.0$ over the sampling time steps. The $T$ value is consistent with the default setting in Stable Diffusion, which is $1000$.
Appendix B A more intuitive explanation of the second stage.
The detailed implementation of MVRS is presented in Alg.2. After completing the first stage of texture generation for the entire room, an initial texture $\mathcal{T}_{MVIS}$ is obtained. However, due to occlusion issues, the initial texture contains numerous untextured black regions. In MVRS, for the n-th viewpoint, the regions with an initial texture are denoted as $P^{n}$, and the corresponding rendered image is represented as $\mathcal{I}_{MVIS}^{n}$. To address the occlusion problem, we perform separate MVRS operations for each instance in the room. After completing the MVRS operations for all instances, we take the union of the texture maps corresponding to each instance to produce the final texture map for the entire scene. This final texture represents the output generated by our proposed framework.
Appendix C Ablation study on the scenond stage
As shown in the Fig.7, using only Stage 1 results in large black areas in the texture caused by occlusion, where textures cannot be generated. By incorporating the second stage , the occlusion issue is resolved while preserving the global style consistency established in the first stage, and the overall texture quality is significantly improved.
Appendix D User Study Details
We conducted the user study using a web-based questionnaire system to compare our method with three baselines from the human perspective. Fig.6 illustrates the interface of our questionnaire system. From the generated results of our method and the baselines, we randomly selected 6 textured scenes created for the same indoor scene using the same textual prompt. In the interface, we first present the textual prompt used for texture generation, followed by rendered images of the textured scene from four different viewpoints. Participants were then asked to evaluate the generated textures across three dimensions: Prompt Fidelity(PF), Texture and Geometry Alignment(TG) and Visual Quality(VQ), with scores ranging from $1$ (low) to $5$ (high). In total, we collected 450 ratings from 25 participants and computed the average score for each method. The result is presented in Tab.2
Figure 6: The interface of the questionnaire system used in user study. We present 4 rendered views from 6 different texturing results to each participant and ask them to rate the scenes across three dimensions.
Appendix E More visual result
E.1 Different style for same room
We present more indoor scene texturing results of our method for one room from 3D-Front in Fig. 8.
E.2 Visual result on ScanNet++ dataset
We are providing multi-view visual results of texturing scenes that from ScanNet++ dataset[47]. It is worth noting that ScanNet++ features more complex layouts compared to 3D-FRONT, as it is a real-world scan dataset. Visual results are shown in Fig.9
E.3 Example of the intermediate denoising process
During the MVIS process of generating multiple consistent images, intermediate results emerge from the multi-view sampling process, as shown in the Fig.10. During MVIS process, the $x_{0,t}$ images across different viewpoints at different steps are shown in Fig.10
Figure 7: Ablation studies on the stage2. Using only Stage 1 results in large black areas in the texture caused by occlusion, where texture cannot be generated. By incorporating Stage 2, the occlusion issue is resolved while preserving the global style consistency established in Stage 1, and the overall texture quality is significantly improved.
Figure 8: Synthesized texture for 3D-FRONT scenes Our method generates various texture for the same input scene by utilizing the prompt template: ’a ⟨STYLE⟩ living room’ with 6 distinct styles for texture creation.
Figure 9: Synthesized texture for ScanNet++ scenes Our method generates similar style texture for different input scene by utilizing the prompt template: ‘a modern style ⟨TYPE⟩’ with 2 distinct room types for texture creation.
Figure 10: Example of MVIS intermedia result An example of the intermediate of Multi-View Integrated Sampling process, showcasing the images $x_{0,t}$ across different viewpoints at different steps.
Algorithm 2 Multi-view Integrated Repaint Sampling of One Instance
  Input: Mesh $\mathcal{M}$, Text $y$, Cameras $\{C^{1},\dots,C^{N}\}$, Textured mask ${\{P^{1},\dots,P^{N}\}}$, Stage1 generated texture $\mathcal{T}_{MVIS}$
  Parameters: DDPM noise schedule $\{\sigma_{t}\}_{t=T}^{0}$
  Initialization:
  $\mathcal{T}=0$
  $\{x_{T}^{n}\}_{n=1}^{N}\sim\{\mathcal{N}({0},{I})\}$ Unpainted noisy latents
  $\{\mathcal{I}_{MVIS}^{n}\leftarrow\mathcal{R}(\mathcal{T}_{MVIS},\mathcal{M},c^{n})\}_{n=1}^{N}$ Painted area is colorful, unpainted area is black
  $\{x_{MVIS}^{n}\leftarrow\mathcal{E}(\mathcal{I}_{MVIS}^{n})\}_{n=1}^{N}$ Painted noise-free latents
  $\triangleright$ Early steps (conditioning on painted area):
  for $t\in\{T\dots T_{end1}\}$ do
     for $n\in\{1\dots N\}$ do
        $\epsilon_{n}\sim\mathcal{N}(\mathrm{0},\boldsymbol{I})$
        $\epsilon_{t}^{n}\leftarrow\epsilon_{\theta}(x_{t}^{n},y,d_{n},t)$
        $\mu_{t-1}^{n}\leftarrow\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}x_{0,t}^{n}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{i}}x_{t}^{n}$
        $x_{t-1}^{n}\leftarrow\mu_{t-1}^{n}+\sigma_{t}\epsilon_{n}$ Unpainted latent with t-1 level noise
        $\tilde{x}_{t-1}^{n}\leftarrow\sqrt{\bar{\alpha}_{t}}x_{MVIS}^{n}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{n}$ Painted latent with t-1 level noise
        $x_{t-1}^{n}\leftarrow\tilde{x}_{t-1}^{n}\odot P^{n}+x_{t-1}^{n}\odot(1-P^{n})$ Mask combine
     end for
  end for
  $\triangleright$ Middle steps (conditioning on painted area):
  for $t\in\{T_{end1}\dots T_{end2}\}$ do
     for $n\in\{1\dots N\}$ do
        $\epsilon_{n}\sim\mathcal{N}(\mathrm{0},\boldsymbol{I})$
        $\epsilon_{t}^{n}\leftarrow\epsilon_{\theta}(x_{t}^{n},y,d_{n},t)$
        $x_{0,t}^{n}\leftarrow\dfrac{x_{t}^{n}-\sqrt{1-\bar{\alpha}_{t}}\bar{\epsilon}_{t}^{n}}{\sqrt{\alpha_{t}}}$ Predicted noise-free image at timestep t
        $\mathcal{I}_{t}^{n}\leftarrow\mathcal{D}(x_{0,t}^{n})$
        $\mathcal{T}^{n}\leftarrow\mathcal{R}^{-1}(\mathcal{I}_{t}^{n},\mathcal{M},C^{n})$ Project view-specific noise-free images to view-specific texture maps
     end for
     $\mathcal{T}=dynamic\_merge(\{\mathcal{T}^{n}\}_{n=1}^{N})$ Merge view-specific texture maps to a global-consistent texture
     for $n\in\{1\dots N\}$ do
        $\tilde{x}_{0,t}^{n}\leftarrow\mathcal{E}(\mathcal{R}(\mathcal{T},\mathcal{M},C^{n}))$ Global-consistent noise-free unpainted area latent at timestep t
        $x_{t-1}^{n}\leftarrow\mu_{t-1}^{n}+\sigma_{t}\epsilon_{n}$ Unpainted latent with t-1 level noise
        $\tilde{x}_{t-1}^{n}\leftarrow\sqrt{\bar{\alpha}_{i}}x_{MVIS}^{n}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{n}$ Painted latent with t-1 level noise
        $x_{t-1}^{n}\leftarrow\tilde{x}_{t-1}^{n}\odot P^{n}+x_{t-1}^{n}\odot(1-P^{n})$ Mask combine
     end for
  end for
  $\triangleright$ Latter steps (not conditioning on painted area):
  for $t\in\{T_{end2}\dots 0\}$ do
     $\mathcal{T}\leftarrow\textbf{MVIS}(\mathcal{M},y,{\{c^{n}\}_{n=1}^{N}},{\{x_{t}^{n}\}_{n=1}^{N}})$
  end for
  $\mathcal{T}_{MVRS}\leftarrow\mathcal{T}$
  return  Texture map $\mathcal{T}_{MVRS}$